OpenStudy (anonymous):

solve pls: 3x+7y=12 -5x+8y=39

OpenStudy (godlovesme):

System of equation?

OpenStudy (anonymous):

i will help hon, give me a sec.

OpenStudy (anonymous):

what is the value of x in the ordered pair that is the solution of the system.....sorry..:) 3x+7y=12 -5+8y=39

OpenStudy (anonymous):

x = 5/3 y = 11/2

OpenStudy (anonymous):

medal or perhaps fan plz?!?!?!

OpenStudy (anonymous):

HELLO, anyone home

OpenStudy (anonymous):

eh are you trying to find the slope or something?

OpenStudy (anonymous):

huh? what is tat supposed to mean

OpenStudy (anonymous):

weirdo

OpenStudy (k_lynn):

@Saisuke<3 was asking if we were supposed to find the slope. Slope is \[y=mx+b\]